OpenStudy (kodak.kiid):

write 22 1/10 as a decimal number?

OpenStudy (jdcap):

\[22\frac{ 1 }{ 10 }=\frac{ 22\cdot10+1 }{ 10 }=\frac{ 221 }{ 10 }\]

OpenStudy (qwertty123):

Hello there! To get a decimal approximation of the fraction, simply divide the fraction's numerator by its denominator. 1 / 10 = ?

OpenStudy (kodak.kiid):

0.1 @Qwertty123

OpenStudy (qwertty123):

Correct! Now add 0.1 to the whole number portion of the mixed number. 0.1 + 22 =?

OpenStudy (jdcap):

NO, you're given a mixed number, not a fraction. It is what I did.

OpenStudy (kodak.kiid):

22.1 @Qwertty123

OpenStudy (jdcap):

Right.. I missed the last step 221/10 = 22.1

OpenStudy (jdcap):

Well... my process has an alternative: \[22\frac{ 1 }{ 10 }=22+\frac{ 1 }{ 10 }=22+0.1=22.1\]

OpenStudy (qwertty123):

Correct! ^.^ Nice Job

OpenStudy (kodak.kiid):

no its 22.05 @Qwertty123

OpenStudy (qwertty123):

0.1 + 22 = 22.1

OpenStudy (kodak.kiid):

oh nvm your right LOL @Qwertty123

OpenStudy (qwertty123):

^.^